19 JANUARY 2000
Starboard Observer: Liz
08:13 Alvin in the water.
08:16 Alvin cleared to descend.
08:33 500 meters.
08:54 1,050 meters.
09:29 Alvin on the bottom; 1,999 meters; 2526X,
09:40 Arrive at Rebeccas Roost; 2616X, 1232Y.
09:45 Surveyed flange with beehive structure on edge.
Active venting, flange surrounded by Riftia. Depth
= 1,987 meters.
09:53 High-temperature probe of beehive. Top = 193°
285° C. Middle = 23° 50° C. Base
= 2° C.
10:00 Began electrochemistry profile of beehive structure.
10:03 Sipper power up. Head registered just fine.
10:30 Some error messages on DLK computer. Called
the surface for advice.
10:33 Sipper 1 fired midway up beehive.
10:35 Powered down DLK, waited, powered it back up,
and after 3 minutes, powered up laptop computer. Everything
10:36 Resumed scanning.
11:13 Pump engaged, Sipper 2 fired outer mid-section
11:27 Pump engaged, Sipper 3 fired 2" above orifice.
11:33 Pump engaged, Sipper 4 fired at base of structure.
11:42 Began sampling sulfide. Though manipulator
arm was in a good position, the sulfide sample fell from
flange. Some pieces recovered.
11:56 High-temperature survey of resulting large
orifice. Gas-tight fired with snorkel in orifice.
12:01 Placed Rebeccas Roost marker
at sampling site.
12:12 Identified new beehive structure; 2613X, 1231.3Y.
Depth = 1,982 meters.
12:21 High-temperature probe profile of new vase-shaped
beehive structure. Top = 66° C, Middle = 80° C.
Structure has large base with skinny chimney on top. Active
12:25 Electrochem. analysis of structure.
12:26 Pump engaged, Sipper 5 fired at top of structure.
12:27 Pump engaged, Sipper 6 fired at mid-section
12:35 Pump engaged, Sipper 7 fired at base of structure.
12:37 Attempt sampling of structure. Beehive fell,
but some pieces recovered.
12:42 Begin transit to new site.
12:58 Triggered aft Niskin off site on way to 2378X,
847Y for coring.
13:17 Arrived at new site; 2381.9X, 832.9Y. Depth
= 1,999.6 meters
13:18 Surveyed bacterial mat with the high-temperature
probe. Surface = 2.5° C, 8" = 95° C. This is a surprising
change in temperature in such a shallow depth. Temperature
monitored as probe removed from sediment.
13:20 Surveyed outside bacterial mat with high-temperature
probe. Surface = 2° C, 3" = 10° C, 8" = 49°
C. Note lower temperature profile outside bacterial mat.
13:22 Began electrochemistry over bacterial mat.
Completed scans to 143. Surface temperature read 20°
50° C on DLK thermistor.
13:27 Began taking cores. Core 3 taken first inside
bacterial mat, Core 5 taken next in bacterial mat, Core
1 taken third outside mat, and Alvin core 2 taken
outside of mat.
13:33 Pump engaged, Sipper 8 fired over mat. Electrochemistry
conducted on same water sample in flow cell.
13:50 Found area of diffuse flow for Niskin, but
too much sediment was stirred up by sub to take water sample.
13:52 Identified group of large crabs. Attempted
to collect one crab with manipulator arm. Collection successful!
13:55 In transit to find a flange with diffuse flow.
Found a great flange at 2371.0X, 822.2Y. Temp = 300°
13:59 Fired two gas-tight syringes under flange.
Excellent video!! Flange is topped with Riftia.
14:12 Second Niskin fired in shimmering water 2357X,
14:15 Surveyed area for possible flange sample, found
14:45 Still surveying area for flange, or tubeworms
surrounded by sediment.
14:50 Identified area of Riftia with sediment.
Alvin core 4 taken, but bad navigation information,
so no XY coordinates for sample.
14:55 Discussion between scientists and pilot regarding
best use of remaining bottom time. Decide to look for sulfide
15:15 On way to look for sulfide, found clam bed.
Decided to collect clams with manipulator arm. Collected
approximately 15 clams. Navigation still bad, no coordinates.
15:25 Alvin cleared to surface.